OpenStudy (anonymous):

Which expression describes the rule of this sequence? 3,6,9,12,15,18,…..? A 3+ (n-1) B 3n C 3+ (3xn) D 3+n

OpenStudy (phi):

do you know that you label each number with its position, starting with 1 1 2 3 4... 3 6 9 12... we call the label "n" so 3 has label 1 and 6 has label 2 if you use rule A : 3+(n-1) and try it for n=2, you should get 6 if it is the correct rule 3+ (2-1)= 3+1= 4. But 4 is not 6. so A is not the right answer. Can you try the other rules and find the correct one?